OpenStudy (anonymous):

All of the following expressions have the same value, when x = -2 and y = 4, except _____. x 2y (-2) 2y -4x 2 -2xy Please help me:(

Directrix (directrix):

Evaluate each of the options for the values x = -2 and y = 4. For the first one, x 2y = (x) * (2) * (y) = (-2) * 2 * (4) = -16. Evaluate (-2) 2y for x = -2 and y = 4 and see if you get -16 or some other product. Continue the process with the remaining two options to find which one is not the same as the others when x = -2 and y = 4.